OpenStudy (anonymous):

solve 7sqr28+3sqr63. solve

OpenStudy (anonymous):

\[7\sqrt{7*2*2}+3\sqrt{7*3*3}\] \[14\sqrt{7}+9\sqrt{7}\] \[23\sqrt{7}\]

OpenStudy (anonymous):

he made a mistake

OpenStudy (anonymous):

I did make a mistake, mathematicalpro is correct. I am deleting my original post. Sorry about that.

OpenStudy (anonymous):

its okay